OpenStudy (anonymous):

Find the distance between the x and y intercepts of 7x-21y=8. Please show your work.

OpenStudy (anonymous):

7x-21y=8. when x=0, 7(0)-21y=8. y=-8/21, the point is (0,-8/21) when y=0, 7x-21(0)=8. 7x=8 x=8/7, the point is (8/7,0) the distance between the point(8/7,0) and (0,-8/21) is d=sqrt[(0-8/7)^2 -(-8/21-0)^2]=1.0775 ans